# Constructing Safety Cases for AI Systems: A Reusable Template Framework

### v1

@@ 3.5 Data synthesis and analysis
We first extracted all relevant information together with the RQ-specific items related to CAE structures and any proposed safety-case templates or patterns. Based on these data, we performed a mixed quantitative and qualitative synthesis aligned with our two research questions (RQ1 and RQ2). On the quantitative side, we used simple descriptive statistics to identify patterns and trends across the selected studies on AI safety cases and structured argumentation.
We summarized (i) the distribution of safety claim types reported in the literature (e.g., absolute, comparative/marginal, constrained/envelope, capability-limited claimes), (ii) the argument strategies used to justify claims (e.g., layered arguments, barrier arguments, comparative arguments), and (iii) the evidence categories employed to support claims and arguments, such as testing, user studies,historical data, and expert assessment.
We also conducted a thematic synthesis of the extracted CAE-related data. We grouped similar concepts and approaches across studies to derive higher-level themes concerning (i) how prior work classifies or organizes claims, arguments, and evidence for AI systems, leading to a structured CAE taxonomy and framework for AI safety cases, (ii) how CAE elements are combined in practice (e.g., common argument structures or evidence bundles that can be expressed as reusable safety-case templates or patterns for AI systems), and (iii) observations such as prevalent reliance on particular argument styles, dominance of certain evidence types, gaps in post-deployment or uncertainty-focused evidence, and examples of innovative or effective CAE practices.
The derived themes were iteratively refined against the primary studies to ensure that the resulting CAE taxonomy and associated templates and patterns accurately reflected the underlying evidence.
@@ 3.6 Summary information
To provide an overview of the publication trends within our SLR dataset, Figure 6 summarises the temporal and typological distribution of the 112 selected studies. Figure 6(a) shows the number of selected papers by publication year; prior to 2020 the number of relevant works remains comparatively low and stable, with only modest growth between 2021 and 2023, but there is a sharp increase from 2024 onwards. Figure 6(b) presents the distribution of article types and indicates that most studies appear as journal articles (40 out of 112, 35.71%), arXiv preprints (35 out of 112, 31.25%) and conference papers (31 out of 112, 27.67%).
(a) Number of selected papers by year; publications show a sharp increase in 2024 and 2025.
@@
(b) Number of selected papers by type; majority are journal, conference and recently posted arXiv papers.
Figure 6: The distribution of the selected articles in this study by published year and article types.
@@
Table 4 organises the CAE taxonomy into its Claim, Argument, and Evidence categories, each further divided into subcategories. For this mapping, in addition to the 112 papers selected through the SLR process, we also incorporated seven supplementary papers identified through targeted search. These papers were not retrieved through the automated search or snowballing phases for two reasons. First, several of them use terminology that predates contemporary CAE- or AI-specific safety-case vocabulary, and second, some fall outside the indexing scope of the chosen databases. Nevertheless, they represent influential or foundational works that provide important conceptual grounding for safety-case thinking, particularly in relation to argument structures, inability-based reasoning, and safety-case reuse.
@@
In Table 4, the "Core reference" column records studies that explicitly instantiate a given subcategory in our analysis. For instance, studies S2, S7, and S8 state assertion-based claims as direct safety guarantees. Similarly, risk-based argumentation is represented where authors substantively engage hazards and mitigations such as S18’s barrier-based reasoning for reducing exfiltration risk. The evidence layer captures explicit support. The empirical entries reflect reported evaluation artefacts such as red-team outcomes (e.g., S18 and S27). The formal-methods evidence is reserved for studies that provide proof-style assurance (e.g., S20) as part of a safety case. This treatment aligns with the presentation of defining categories with concrete, method-relevant examples.
@@
The "Inferred mapping" column extends coverage by recording studies whose content functionally aligns with a subcategory even when the authors do not label it as such. This improves completeness but also introduces a methodological trade-off, i.e., implicit assignments can blur the boundary between observed CAE framing and inferred alignment. This increases susceptibility to interpretive bias unless the inference rules are explicit and consistently applied such as for S73 within casual and explanatory argument. The following example illustrates both the benefit and the risk considering example studies:
@@
Example Mapping study S112 under risk-based argument highlights a substantive critique that safety cases should priorities risk reduction over compliance. Similarly, adding S64 under operational/field evidence reflects an evidential stance that assumes runtime monitoring artifacts. However, these inferences should be read as analytical triangulation, not a re-labelling of author intent. We believe the table is strongest when implicit links are justified via clear criteria e.g., required artifact types, explicit causal reasoning, or risk-management constructs and reported as a validity mitigation rather than a substitute for explicit CAE usage in the studies.
4 A Framework of Reusable AI Safety-Case Templates
@@ 4.1 Rationale for reusable templates
The emerging diversity of AI systems, from narrow domain models to large-scale generative models, demands a systematic yet adaptable way to structure safety justification. In traditional engineering, safety cases are often bespoke, created for each project through expert judgement and iterative review. For AI, this approach does not scale. Model behaviours evolve too quickly, evidence sources are heterogeneous, and assurance must often be performed by teams without formal safety-case experience.
A reusable, pattern-based framework can reduce subjectivity and promote consistency, while remaining flexible enough to handle uncertainty, discovery, and change. Following [bommasani2021opportunities] and adapting their template-based assurance logic for autonomous systems, the proposed framework defines reusable safety-case templates as standardised, semi-structured argument skeletons. Each template combines a claim pattern, a corresponding argument structure, and a set of evidence families appropriate to AI systems. These templates can be specialised or composed to construct end-to-end safety cases.
@@ 4.2 Overview of the architecture
The framework follows the classical CAE triad but adapts each element to AI-specific realities (Table 5).
@@
Evidence Test results, analyses, and certifications Empirical evaluations, red-teaming, interpretability analyses, governance artefacts, and risk-model outputs
@@
Figure 7 shows the full CAE taxonomy for AI systems, which forms the structural foundation of the proposed reusable safety-case template. The taxonomy integrates AI-specific extensions into the classical CAE model, distinguishing how safety propositions (claims), reasoning structures (arguments), and supporting artefacts (evidence) interrelate across assurance layers.
@@ 4.3 Claim
Constrained claims express that safety is conditional on defined operational limits such as data regime, modality, or environmental boundary. They ensure that safety justification remains valid only within those bounds.
  • Constrained claim: Expresses that safety is bounded by particular operational conditions, data regimes, or user contexts. Context and constraint are treated jointly, as they are inseparable in AI (e.g., domain, modality, data source, tool-access limits).
@@
This type of claim states that safety depends on capability boundaries, either control-based (design choices, such as tool or network disablement), or intrinsic (assumed AI system inability or refusal behaviour).
@@
    • Actuation-limited aligns with prior safety research advocating containment and human-in-command architectures, which restrict AI systems from issuing direct actuation commands to prevent uncontrolled or unsafe physical actions.
@@
    • Access-limited asserts that the system cannot access unauthorised networks, data repositories, or sensitive resources. It limits data-exfiltration and privacy risks.
@@
Each claim type implies specific expectations for argument structure and evidence. For instance, contextual claims require evidence of boundary enforcement and capability-limited claims require mechanism inspection or adversarial validation.
@@ 4.4 Argument
Arguments provide the connective reasoning that links claims to supporting evidence. AI safety cases require argument structures that can integrate heterogeneous reasoning modes and adapt as new evidence emerges. The proposed taxonomy defines two orthogonal dimensions including function and logical form, which can be combined modularly.
@@ 4.4.1 Argument taxonomy
Comparative arguments use inductive and statistical reasoning for comparative argument and abductive reasoning for Comparison-based argument. They are supported by benchmarking, controlled evaluations, and expert-review evidence.
@@
  • Guideline-based argument demonstrates that the system complies with an established safety or risk-management framework and infers acceptability from that compliance. It argues indirectly that adherence to recognised safety processes implies acceptable risk.
@@
While typical reasoning for Threshold-based argument is statistical logic, deductive reasoning is used for Guideline-based argument.
@@
    Example: “We justify the safety of System X by explaining and eliminating the previously observed risk. We observed [O1–O3] and evaluated competing hypotheses [H0…Hn] using [D1…Dk] (ablations, audits, counterfactuals). Mitigation M removes the hazard and M was deployed."
@@
    Example: “System X is capable to cause serious harm, but it is restricted from using those capabilities; The AI system will not be modified such that it can use its capabilities to cause serious harm; The AI system’s model weights will not be stolen."
@@
    Example: “System X is not capable to cause serious harm, regardless of how it is modified or used; Safety-alignment training enforces refusal to generate disallowed instructions across tested scenarios.”
@@
Normative/conformance arguments show adherence to established frameworks, governance standards, or dependability engineering principles that are widely recognised as proxies for safety assurance. This argument type demonstrates that system safety is justified through compliance with normative expectations rather than new empirical discovery. This argument includes two types of argument: Guideline-based argument and Dependability argument. Although these argument types also appear under other functional categories (e.g., risk-based or causal/explanatory), the taxonomy is not exclusive: certain reasoning forms can legitimately contribute to multiple assurance dimensions depending on the intent and context of their use. In this category, the emphasis is on institutional acceptability and auditability rather than on demonstrating new technical risk reduction. Compliance with recognised frameworks or dependability doctrines functions as a proxy for acceptable safety within established governance regimes.
  • Guideline-based argument demonstrates that the system complies with an established safety or risk-management framework and infers acceptability from that compliance. It argues indirectly that adherence to recognised safety processes implies acceptable risk.
@@ 4.4.2 Reasoning logic
It provides rigour and traceability and supports formal verification or policy-based reasoning [basir2009deriving]. Yet, it requires stable premises and rarely sufficient for emergent or probabilistic behaviours.
@@
"A new developed “AI-based document evaluation system” has four safety layers (policy, testing, controls, and monitoring). Each layer meets its defined acceptance criteria, and all interfaces between layers are verified.
@@
  • Policy layer: The system have documented risk classification and approval before deployment.
@@
While this logic captures real-world performance and adapts to continual evaluation, confidence depends on representativeness and coverage of the test conditions and induction cannot guarantee future behaviour [leake1996case].
@@
"Scenario 1: System X is an updated LLM model of System Y and Z (previous version of the model), both of which have been used safely in production for over a year. Since the lasted version of model meets or exceeds their safety benchmarks, it is inferred that it is also safe."
"Scenario 2: System X is a new LLM. After comparison with widely used LLMs (e.g., ChatGPT and Gemini), its performance on safety benchmarks is equal to or better than these reference systems. Therefore, System X is considered safe and acceptable."
@@
It enables learning-oriented assurance and also turns unexplained anomalies into explainable risks with documented hypotheses [burton2022causal]. Its limitations include reliance on expert judgment and explanation plausibility, and susceptibility to confirmation bias without cross-validation.
@@
"AI-based risk management system kept crashing after long use. Engineers hypothesised it was due to memory leaks (H1), corrupted user data (H2), and overheating (H3). After testing/observation, confirmed memory leaks was the root cause (i.e., updated the memory management module, and the crashes stopped)."
@@
"Process analysis AI system has been newly developed. The organisation’s risk threshold for AI system is 0.25. Based on testing and field data, the system’s estimated risk score is 0.18, below the organisation’s approved threshold of 0.25. Since risk score $\leq$ threshold, the system is considered acceptable for deployment."
@@
Meta-assurance, reasoning about the quality of reasoning, becomes essential. It requires showing that each logic is used under its valid conditions, that assumptions are traceable, and that uncertainty is bounded both within and across reasoning modes. For example, each logical form includes an associated pass condition that defines its validity criteria. These pass conditions collectively enable meta-assurance, by allowing the reasoning process itself to be verified for correctness, coherence, and justified application.

### v3

@@ 3.5 Data synthesis and analysis
We first extracted all relevant information, including the RQ-specific items related to CAE structures and any proposed safety-case templates or patterns. Based on these data, we conducted a mixed-methods synthesis aligned with our two research questions (RQ1 and RQ2). On the quantitative side, we used simple descriptive statistics to identify patterns and trends across the selected studies on AI safety cases and structured argumentation.
In addition to descriptive distributions, we conducted an explicit synthesis to characterize the state of the art in AI safety-case construction. This synthesis identifies (i) recurring strengths and weaknesses in how claims are formulated, e.g., conditional vs. unconditional safety, (ii) how arguments establish acceptability, e.g., reliance on comparators without explicit thresholds, (iii) how evidence is operationalized, e.g., evaluation artefacts without stated quality criteria or traceability. We then map these findings to reusable template requirements, including fields, constraints, and update triggers, so that the study directly addresses observed gaps.
We also conducted a thematic synthesis of the extracted CAE-related data. We grouped similar concepts and approaches across studies to derive higher-level themes concerning (i) how prior work classifies or organizes claims, arguments, and evidence for AI systems, leading to a structured CAE taxonomy and template for AI safety cases, (ii) how CAE elements are combined in practice e.g., common argument structures or evidence bundles that can be expressed as reusable safety-case templates or patterns for AI systems, and (iii) observations such as prevalent reliance on particular argument styles, dominance of certain evidence types, gaps in post-deployment or uncertainty-focused evidence, and examples of innovative or effective CAE practices. The derived themes were iteratively refined against the primary studies to ensure that the resulting CAE taxonomy, along with its associated templates and patterns, accurately reflected the underlying evidence.
@@ 3.6 Summary information
To provide an overview of the publication trends within our SLR dataset, Figure 6 summarises the temporal and typological distribution of the 112 selected studies. Figure 6(a) shows the number of selected papers by publication year; prior to 2020, the number of relevant works remains comparatively low and stable, with only modest growth between 2021 and 2023, but there is a sharp increase from 2024 onwards. Figure 6(b) presents the distribution of article types and indicates that most studies appear as journal articles (40 out of 112, 35.71%), arXiv preprints (35 out of 112, 31.25%), and conference papers (31 out of 112, 27.67%).
@@
Figure 6: The distribution of the selected articles in this study by published year and article types. (a) Number of selected papers by year; publications show a sharp increase in 2024 and 2025.(b) Number of selected papers by type; the majority are journal, conference, and recently posted arXiv papers.
3.6.1 Overview of the selected studies
@@
Table 5 organises the CAE taxonomy into its Claim, Argument, and Evidence categories, each further divided into subcategories. For this mapping, in addition to the 112 papers selected through the SLR process, we also incorporated seven supplementary papers identified through a targeted search. These papers were not retrieved through the automated search or snowballing phases for two reasons. First, several of them use terminology that predates contemporary CAE- or AI-specific safety-case vocabulary, and second, some fall outside the scope of indexing by the chosen databases. Nevertheless, they represent influential or foundational works that provide important conceptual grounding for safety-case thinking, particularly regarding argument structures, inability-based reasoning, and safety-case reuse.
@@
In Table 5, the "Core reference" column records studies that explicitly instantiate a given subcategory in our analysis. For instance, studies S2, S7, and S8 state assertion-based claims as direct safety guarantees. Similarly, risk-based argumentation is represented where authors substantively engage hazards and mitigations, such as S18’s barrier-based reasoning to reduce exfiltration risk. The evidence layer captures explicit support. The empirical entries reflect reported evaluation artefacts such as red-team outcomes (e.g., S18 and S27). The formal-methods evidence is reserved for studies that provide proof-style assurance (e.g., S20) as part of a safety case. This treatment aligns with the presentation of defining categories with concrete, method-relevant examples.
@@
The "Inferred mapping" column extends coverage by recording studies whose content functionally aligns with a subcategory even when the authors do not label it as such. This improves completeness but also introduces a methodological trade-off: implicit assignments can blur the boundary between the observed CAE framing and the inferred alignment. This increases susceptibility to interpretive bias unless the inference rules are explicit and consistently applied, as in S73 within causal and explanatory arguments. The following example illustrates both the benefit and the risk, considering example studies:
@@
Example The mapping study S112, under a risk-based argument, highlights a substantive critique: safety cases should prioritize risk reduction over compliance. Similarly, adding S64 under operational/field evidence reflects an evidential stance that assumes the presence of runtime monitoring artifacts. However, these inferences should be read as analytical triangulation, not a re-labelling of author intent. We believe the table is strongest when implicit links are justified via clear criteria, e.g., required artifact types, explicit causal reasoning, or risk-management constructs, and reported as a validity mitigation rather than a substitute for explicit CAE usage in the studies.
3.6.2 Key insights into AI safety case construction
Across the 112 analyzed studies, AI safety cases most commonly take the form of conditional claims supported by empirical evidence, often because ‘ground truth’ safety is unavailable in real deployments. However, we observe three recurring shortcomings.
  • Acceptability criteria are underspecified. Comparative or marginal claims frequently rely on a baseline (no worse than thresholds) without defining the comparator system’s assurance status or formalizing failure budgets and quantitative thresholds.
  • Evidence quality criteria and traceability are inconsistently stated. While red-teaming, stress testing, and audit artefacts appear widely, many safety-case variants do not specify evidence independence, coverage, recency, reproducibility, or linkage, which limits auditability under continuous updates.
  • Post-deployment assurance remains fragmented. A smaller subset of studies ($\leq 10\%$) link claims to operational monitoring and governance actions, which becomes increasingly necessary as models are updated, tool access changes, or deployment contexts drift.
Overall, these three major findings motivate our template requirements, i.e., each template instance must explicitly declare comparators/thresholds (when used), evidence quality criteria, and update triggers that require re-validation.
The reviewed literature indicates that AI safety case construction remains at a formative stage. While the CAE structure is widely adopted in form, its operationalisation varies significantly in claim definition, evidential adequacy, and lifecycle integration. The absence of reusable abstraction layers and formalised update logic limits composability and long-term auditability. This structural diagnosis motivates the need for systematic taxonomies and reusable template patterns to stabilise AI safety case engineering across domains.
4 Reusable AI Safety-Case Templates
@@ 4.1 Rationale for reusable templates
The emerging diversity of AI systems, from narrow-domain models to large-scale generative models, demands a systematic yet adaptable approach to structuring safety justification. In traditional engineering, safety cases are often bespoke, created for each project through expert judgement and iterative review. For AI, this approach does not scale. Model behaviours evolve too quickly, evidence sources are heterogeneous, and assurance must often be performed by teams without formal safety-case experience.
A reusable, pattern-based approach can reduce subjectivity and promote consistency, while remaining flexible enough to handle uncertainty, discovery, and change. Following [bommasani2021opportunities] and adapting their template-based assurance logic for autonomous systems, the proposed solution defines reusable safety-case templates as standardised, semi-structured argument skeletons. Each template combines a claim pattern, a corresponding argument structure, and a set of evidence families appropriate to AI systems. These templates can be specialised or composed to construct end-to-end safety cases.
@@ 4.2 Overview of the architecture
The structure follows the classical CAE triad but adapts each element to AI-specific realities (Table 6).
@@
Evidence Test results, analyses, certifications Empirical evaluations, red-teaming, interpretability analyses, governance artefacts, and risk-model outputs
@@
Figure 7 shows the full CAE taxonomy for AI systems, which serves as the structural foundation for the proposed reusable safety-case template. The taxonomy integrates AI-specific extensions into the classical CAE model, distinguishing how safety propositions (claims), reasoning structures (arguments), and supporting artefacts (evidence) interrelate across assurance layers.
Figure 7: The CAE Taxonomy for AI systems.
@@ 4.3 Claim
Constrained claims state that safety is conditional on defined operational limits, such as data regimes, modalities, or environmental boundaries. They ensure that safety justification remains valid only within those bounds.
  • Constrained claim: Expresses that safety is bounded by particular operational conditions, data regimes, or user contexts. Context and constraints are treated jointly, as they are inseparable in AI (e.g., domain, modality, data source, tool access limits).
@@
This type of claim states that safety depends on capability boundaries, either control-based (design choices, such as tool or network disablement) or intrinsic (assumed AI system inability or refusal behaviour).
@@
    • Actuation-limited aligns with prior safety research advocating containment and human-in-command architectures, which restricts AI systems from issuing direct actuation commands to prevent uncontrolled or unsafe physical actions.
@@
    • Access-limited asserts that the system cannot access unauthorised networks, data repositories, or sensitive resources. It limits data exfiltration and privacy risks.
@@
Each claim type implies specific expectations for argument structure and evidence. For instance, contextual claims require evidence of boundary enforcement, whereas capability-limited claims require inspection of mechanisms or adversarial validation.
@@ 4.4 Argument
Arguments provide the connective reasoning that links claims to supporting evidence. AI safety cases require argument structures that can integrate heterogeneous reasoning modes and adapt as new evidence emerges. The proposed taxonomy defines two orthogonal dimensions, including function and logical form, which can be combined modularly.
@@ 4.4.1 Argument taxonomy
Comparative arguments use inductive, statistical, and abductive reasoning. They are supported by benchmarking, controlled evaluations, and evidence from expert reviews.
@@
  • Guideline-based argument demonstrates that the system complies with an established safety or risk-management framework and infers acceptability from that compliance. It indirectly argues that adherence to recognised safety processes implies an acceptable level of risk.
@@
While the typical rationale for a threshold-based argument is statistical reasoning, a guideline-based argument relies on deductive reasoning.
@@
    Example: “We justify the safety of System X by explaining and eliminating the previously observed risk. We observed [O1–O3] and evaluated competing hypotheses [H0…Hn] using [D1…Dk] (ablations, audits, counterfactuals). Mitigation M removes the hazard, and M was deployed."
@@
    Example: “System X is capable of causing serious harm, but it is restricted from using those capabilities; The AI system will not be modified such that it can use its capabilities to cause serious harm; The AI system’s model weights will not be stolen."
@@
    Example: “System X is not capable of causing serious harm, regardless of how it is modified or used; Safety-alignment training enforces refusal to generate disallowed instructions across tested scenarios.”
@@
Normative/conformance arguments demonstrate adherence to established frameworks, governance standards, or dependability engineering principles widely recognised as proxies for safety assurance. This argument type demonstrates that system safety is justified through compliance with normative expectations rather than new empirical discovery. This argument includes two types: Guideline-based argument and Dependability argument. Although these argument types also appear under other functional categories (e.g., risk-based or causal/explanatory), the taxonomy is not exclusive: certain reasoning forms can legitimately contribute to multiple assurance dimensions depending on the intent and context of their use. In this category, the emphasis is on institutional acceptability and auditability rather than on demonstrating reductions in new technical risk. Compliance with recognised frameworks or dependability doctrines functions as a proxy for acceptable safety within established governance regimes.
  • Guideline-based argument demonstrates that the system complies with an established safety or risk-management framework and infers acceptability from that compliance. It indirectly argues that adherence to recognised safety processes implies an acceptable level of risk.
@@ 4.4.2 Reasoning logic
It provides rigour and traceability and supports formal verification or policy-based reasoning [basir2009deriving]. Yet, it requires stable premises and is rarely sufficient for emergent or probabilistic behaviours.
@@
"A newly developed “AI-based document evaluation system” has four safety layers (policy, testing, controls, and monitoring). Each layer meets its defined acceptance criteria, and all interfaces between layers are verified.
@@
  • Policy layer: The system has documented risk classification and approval before deployment.
@@
While this logic captures real-world performance and adapts to continual evaluation, confidence depends on representativeness and coverage of the test conditions, and induction cannot guarantee future behaviour [leake1996case].
@@
"Scenario 1: System X is an updated LLM model of System Y and Z (previous version of the model), both of which have been used safely in production for over a year. Since the latest version of the model meets or exceeds their safety benchmarks, it is inferred that it is also safe."
"Scenario 2: System X is a new LLM. After comparison with widely used LLMs (e.g., ChatGPT and Gemini), its performance on safety benchmarks is equal to or better than that of these LLMs. Therefore, System X is considered safe and acceptable."
@@
It enables learning-oriented assurance and transforms unexplained anomalies into explainable risks with documented hypotheses [burton2022causal]. Its limitations include reliance on expert judgment and the plausibility of explanations, as well as susceptibility to confirmation bias without cross-validation.
@@
"AI-based risk management system kept crashing after long use. Engineers hypothesised it was due to memory leaks (H1), corrupted user data (H2), and overheating (H3). After testing/observation, confirmed memory leaks were the root cause (i.e., updated the memory management module, and the crashes stopped)."
@@
"Process analysis AI system has been newly developed. The organisation’s risk threshold for the AI system is 0.25. Based on testing and field data, the system’s estimated risk score is 0.18, below the organisation’s approved threshold of 0.25. Since risk score $\leq$ threshold, the system is considered acceptable for deployment."
@@
Meta-assurance, reasoning about the quality of reasoning, becomes essential. It requires showing that each logic is used under its valid conditions, that assumptions are traceable, and that uncertainty is bounded both within and across reasoning modes. For example, each logical form includes an associated pass condition that defines its validity criteria. These pass conditions collectively enable meta-assurance by allowing the reasoning process itself to be verified for correctness, coherence, and justified application.